A polygon has n sides. Two of its angles are right angles. Each of the remaining angles equal to 150 degrees. calculate n. Please help! =)
Answer by Alan3354(69443)

A polygon has n sides. Two of its angles are right angles. Each of the remaining angles equal to 150 degrees. calculate n.
--------------------
n = # of sides
Sum of Int angles = 180*(n-2) degs
180*(n-2) = 90 + 90 + (n-2)*150
180n - 360 = 180 + 150n - 300 = 150n - 120
30n = 240
n = 8